OpenStudy (anonymous):

find f^-1 ' (0) for f(x)= cosx on [0, pi/2]

OpenStudy (anonymous):

since f(pi/2) = cos(pi/2) = 0, then f^(-1)(0) = pi/2 = arccos(0)